• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name dalton

    Log entry time 16:37:51 on October 24, 2009

    Entry number 296461

    This entry is a followup to: 296312

    keyword=Compton laser polarisation fixed

    From the quarter wave plate (the one in the analysis station) scan it was
    possible to determine that the problem was with the quarter wave plate
    before the cavity (the one that changes the helicity of the photon beam).
    Presumably it was over or under rotating when changing from one helicity
    to another, leaving a residual component of linear polarisation in the beam.

    The most bizarre thing of all: an epics reboot fixed the problem. A
    further QWP scan poduced the familiar output.

    Figure 1